OpenStudy (anonymous):

what is the result of isolating x^2 in the equation below 10x^2-36y^2=100

OpenStudy (aum):

10x^2-36y^2=100 To isolate x^2, first add 36y^2 to both sides: 10x^2 = 100 + 36y^2 divide both sides by 10: x^2 = 10 + 3.6y^2 or 10 + 36/10 * y^2 = 10 + 18/5 * y^2

OpenStudy (aum):

\[\large x^2 = 10 + \frac{18}{5}y^2 \]